OpenStudy (anonymous):

evaluate each expression when x is 1,2,3, and 4 for 3x?

OpenStudy (mayankdevnani):

1) 3*1=3 2)3*2=6 3)3*3=9 4)3*4=12

OpenStudy (mayankdevnani):

@shivraj ok

OpenStudy (anonymous):

thnks

OpenStudy (mayankdevnani):

welcome

OpenStudy (mayankdevnani):

any confusion